OpenStudy (anonymous):

What is the simplified form of 3/4x+3 + 21/x^2+26x+15 ? someone help please

OpenStudy (pfenn1):

Is your expression\[\left( 3 \over 4x+3 \right)+\left( 21\over x ^{2}+26x+15 \right)\]

OpenStudy (anonymous):

the second equation is supposed to be 21/ 8x^2 +26x+15

OpenStudy (pfenn1):

you mean\[\left( 21 \over 8x ^{2} \right)+26x+15\]

OpenStudy (pfenn1):

or you mean the other was right but was missing the 8?

OpenStudy (anonymous):